OpenStudy (mercury):

A professor gives some students a test with a time limit of 1 hour, then hops on his spacecraft and travels at 9,97*c. When HIS clock says 1 hour has passed, he uses a prototype quantum entanglement device that INSTANTLY sends a signal back to earth. How much time did the students have?

OpenStudy (mercury):

Typing error: velocity is 0,97*c (...duh)

OpenStudy (bugman954):

t0 = t/(1-(v/c)^2)^(1/2) t0 = 1hour/(1-(0.97/1)^2)^(1/2) t0 = 4.113450349hours http://abyss.uoregon.edu/~js/ast122/lectures/lec20.html

OpenStudy (mercury):

Well, not what I was looking for. See, the Lorentz transformation implies that the signal is sent at light speed. But what if the signal was sent simultaneously ( ignoring the apparent break in causality) ?

OpenStudy (bugman954):

@Mercury From your question "quantum entanglement device that INSTANTLY sends a signal" the professor realizes time dilation where his hour is up after 4.113450349hours of testing in the slower frame and he "INSTANTLY" stops the test using his "quantum entanglement device" Maybe you need to brush up on quantum entanglement theory where there is no delay in the change in state of the entangled particles regardless of distance.

OpenStudy (mercury):

Actually, I have no background in qe,:P but I do know that all information is communicated at the speed of light. If you use a minskowski diagram, you can see that communication between the professor and the students has to be through the speed of light. Well, I want to know what would happen if no time would be required for the two systems to communicate an event. If the quantum entanglement theory poses problems, just think of it as an instantaneous communication. Thanks for your time anyway.

OpenStudy (mercury):

After speaking with my professor ( who doesn't yet posses such spaceship 😔), I see that you are correct. So, if the signal were to travel at light speed, the students would notice it at to = 8.2 hours. Thank you